# Centre for Cyber-Physical Systems IISc, Bengaluru, India

### v1

@@ 4.4 Rotating Spacecraft Model
where $\mathsf{x}=[\mathsf{x}_{1},\mathsf{x}_{2},\mathsf{x}_{3}]^{\top}$ denotes angular velocities $\omega_{1},\omega_{2},\omega_{3}$ along the principal axes respectively, $\mathsf{u}=[\mathsf{u}_{1},\mathsf{u}_{2},\mathsf{u}_{3}]^{\top}$ represents the torque input, and with $J_{1}=200,J_{2}=200,J_{3}=100$ denote the principal moments of inertia. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the input set to be bounded within ${\mathbf{W}}=[-1,1]$. Note that the internal input is of 3 dimensions while external input is of a single dimension, i.e., ${\mathbf{U}}\in{\mathbb{R}}^{3}$, but ${\mathbf{W}}\in{\mathbb{R}}$. Also, we consider the model to be unknown. However, we estimate the Lipschitz constants $\mathcal{L}_{x}=1,\mathcal{L}_{u}=0.01$.
The goal is to synthesize a controller to enforce the system to be $\delta$-ISS. To do this, we first fix the training hyper-parameters as $\epsilon=0.0125,\mathcal{L}_{L}=1,\mathcal{L}_{C}=40,k_{1}=0.00001,k_{2}=0.5,k_{3}=0.001,k_{w}=0.01$. So, the Lipschitz constant according to Theorem 3.2 is $3.651$. We fix the structure of $V_{\theta,b}$ as $l_{f}=1,h_{f}^{1}=60$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=40$. The training algorithm converges to obtain $\delta$-ISS-CLF $V_{\theta,b}$ along with $\eta=-0.0460$. Hence, $\eta+\mathcal{L}\epsilon=-0.0460+3.651\times 0.0125=-0.00036$, therefore, using Theorem 3.5, we can guarantee that the obtained $\delta$-ISS-CLF $V_{\theta,b}$ is valid and that the closed-loop system is assured to be incrementally input-to-state stable under the influence of the controller $g_{\bar{\theta},\bar{b}}$.
The successful runs of the algorithm have an average convergence time of 6.5 hours.
@@ 4.5 Discussion
The controller synthesis technique proposed in this work for achieving incremental input-to-state stability in closed-loop systems offers several advantages over existing methods. As the current technique relies on synthesizing the controller using data of the system, hence designing the controller does not require the system to be either fully actuated or affine structure as done in previous conventional techniques [19, 20]. This neural network based controller synthesis is able to make underactuated systems incrementally stable without the knowledge of dynamics. For instance, our neural network-based controller successfully ensures incremental stability in the one-link manipulator and jet-engine model without requiring detailed dynamic equations. The approach is also reliable for the case of nonaffine systems as well, as seen in the scalar example.
Also, from the final case study, one can generate three different control inputs using a single external input, so the closed-loop system is controllable with one single input, thereby simplifying control implementation. Now, as seen from [4], the computation complexity for the abstraction of an incrementally stable system lies in the dimension of the input to the system. As we have made the system incrementally stable with one input only, the computational complexity of the abstraction will reduce drastically once the controller is trained. This is helpful in finding the scalable symbolic abstraction of the system and, thus, will enable the efficient design of another formally verified controller to perform other specifications for the closed-loop system.
@@ 5 Conclusion and Future Work
In this paper, we presented a new training approach for unknown discrete-time systems that simultaneously synthesizes a controller and a verifiably correct incremental input-to-state stable control Lyapunov function, both parameterized as neural networks. The notion of incremental input-to-state state control Lyapunov function was introduced for the very first time. We highlighted the crucial challenges that are faced while solving an ROP to ensure the existence of $\delta$-ISS-CLF and thereby mapped the problem into SOP. Consequently, we presented a validity condition, by virtue of which solving SOP ensures satisfying the corresponding ROP, necessitating the existence of appropriate $\delta$-ISS-CLF. We illustrated four case studies in support of our proposed claims. Our current approach was limited to the framework of unknown discrete time control systems. However, the framework can easily be extended to construct $\delta$-ISS-CLF under the action of the controller for continuous-time systems, which is now in an investigation stage. Moreover, we plan to extend these results to guarantee incremental ISS for the unknown systems in stochastic settings and large-scale interconnected systems so that the proposed results will be further generalized and scalable.
Appendix A. Proof of Theorem 2.8
The proof follows a similar approach of [13, Theorem 8]. Let the closed-loop system admit a $\delta$-ISS-CLF satisfying the conditions of Definition 2.7. The lower bound (i) and inequality (ii) imply: $\forall x,\hat{x}\in{\mathbf{X}},w,\hat{w}\in{\mathbf{W}}$:
$V(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))-V(x,\hat{x})\leq-\alpha(V(x,\hat{x}))+\sigma(|w-\hat{w}|),$ (20)
@@
We denote $id$ as an identity function, i.e. $id(s):=s$ for all $s\in{\mathbb{R}}_{0}^{+}$. Assume $\chi$ be any class $\mathcal{K}_{\infty}$ such that $id-\chi\in\mathcal{K}_{\infty}$. Without loss of generality, we assume $id-\alpha\in\mathcal{K}$ [38, Lemma B.1].
Let ${\mathsf{w}},\hat{{\mathsf{w}}}$ denote two input sequences such that ${\mathsf{w}},\hat{{\mathsf{w}}}:{\mathbb{N}}_{0}\rightarrow{\mathbf{W}}$. Now, define the set :
$\mathbb{S}:$ =$=\{(\epsilon_{1},\epsilon_{2})|\epsilon_{1},\epsilon_{2}\in{\mathbf{X}},V(\epsilon_{1},\epsilon_{2})\leq\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)\}.$ (21)
@@
Proof 5.2
Let $(x,\hat{x})\in\mathbb{S}$. Also, ${\mathsf{w}}(0)=w,\hat{{\mathsf{w}}}(0)=\hat{w}\in{\mathbf{W}}$ be the first elements of ${\mathsf{w}},\hat{{\mathsf{w}}}$ respectively. Note that the set ${\mathbf{X}}$ is compact and hence under the controller $g$, it is control forward invariant. Therefore, the terms $f(x,g(x,w))$ and $f(\hat{x},g(\hat{x},\hat{w}))\in{\mathbf{X}}$.
Then, $V(x,\hat{x})\leq\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)$ and combining with (20), we have:
$V(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))$
$\leq V(x,\hat{x})-\alpha(V(x,\hat{x}))+\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)$
$=(id-\alpha)V(x,\hat{x})+\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)$
$\leq(id-\alpha)\circ\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)+\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)$
$=\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)-\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)+\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)$
$=\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)-(id-\chi)\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert).$ (22)
Since $(id-\chi)\in\mathcal{K}_{\infty}$, hence $-(id-\chi)\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)\leq 0$. So, essentially,
$V(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))\leq\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)$ (23)
which implies that $(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))\in\mathbb{S}$. Hence, we can say that the set $\mathbb{S}$ is forward invariant.
@@
Part-I: For $(x,\hat{x})\in\mathbb{S}$, applying condition (i) of Definition 2.7, we obtain:
$\alpha_{1}(|\mathsf{x}(k,x,g(x,\mathsf{w}(k))-\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k)))|)$
$\leq V(\mathsf{x}(k,x,g(x,\mathsf{w}(k))),\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k))))$
$\leq\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert).$
Therefore,
$|\mathsf{x}(k,x,g(x,\mathsf{w}(k)))-\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k)))|\leq\alpha_{1}^{-1}\circ\alpha^{-1}\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)=:\bar{\gamma}(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert),$ (24)
@@
Part-II: For $(x,\hat{x})\notin\mathbb{S}$, (21) shows that $\chi\circ\alpha(V(x,\hat{x}))>\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)$. Hence, (20) gives
$V(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))-V(x,\hat{x})$
$\leq-\alpha(V(x,\hat{x}))+\chi\circ\alpha(V(x,\hat{x}))$
$\leq(id-\chi)\circ\alpha(V(x,\hat{x})).$ (25)
Since, $id-\chi\in\mathcal{K}_{\infty}$, using the comparison lemma [39, Lemma 4.3], we infer there exists some $\tilde{\beta}\in\mathcal{KL}$ so that,
$\alpha_{1}(|\mathsf{x}(k,x,g(x,\mathsf{w}(k)))-\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k)))|)$
$\leq V(\mathsf{x}(k,x,g(x,\mathsf{w}(k))),\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k))))$
@@
$|\mathsf{x}(k,x,g(x,\mathsf{w}(k)))-\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k)))|$
$\leq\alpha_{1}^{-1}\circ\tilde{\beta}(\alpha_{2}(|x-\hat{x}|),k)=:\bar{\beta}((|x-\hat{x}|),k),$ (26)
@@
Now, combining (24) and (Appendix A. Proof of Theorem 2.8), we obtain
$|\mathsf{x}(k,x,g(x,\mathsf{w}(k)))-\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k)))|$
$\leq\max\{\bar{\beta}((|x-\hat{x}|),k),\bar{\gamma}(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)\}$
$\leq\bar{\beta}((|x-\hat{x}|),k)+\bar{\gamma}(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert).$
@@
[…]
@@
$\quad-V_{\theta,b}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})),f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r})))+V_{\theta,b}(x_{q},x_{r})-V_{\theta,b}(x_{q},x_{r})+k_{3}|x-\hat{x}|^{\gamma_{3}}$
@@
$\leq\eta_{S}^{*}+\big{(}\sqrt{2}\mathcal{L}_{L}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+1)+2(\mathcal{L}_{3}+\mathcal{L}_{w})\big{)}\varepsilon$
$\leq\eta_{S}^{*}+\mathcal{L}\varepsilon\leq 0.$
[…]
$=h(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-h(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))+h(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))-k_{h}h(x)+k_{h}h(x_{q})-k_{h}h(x_{q})$
$\leq\mathcal{L}_{h}(f(x,g_{\bar{\theta},\bar{b}}(x,w))-f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))+\eta_{S}^{*}+k_{h}\mathcal{L}_{h}\varepsilon$
$\leq\mathcal{L}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C})\varepsilon+\eta_{S}^{*}+k_{h}\mathcal{L}_{h}\varepsilon\leq\eta_{S}^{*}+\mathcal{L}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+k_{h})\varepsilon\leq\eta_{S}^{*}+\mathcal{L}\varepsilon\leq 0.$
Therefore, if the condition (10) is satisfied, the neural $\delta$-ISS-CLF will ensure the system is incrementally input-to-state stable under the action of the controller. This completes the proof.

### v2

@@ 4.4 Rotating Spacecraft Model
where $\mathsf{x}=[\mathsf{x}_{1},\mathsf{x}_{2},\mathsf{x}_{3}]^{\top}$ denotes angular velocities $\omega_{1},\omega_{2},\omega_{3}$ along the principal axes respectively, $\mathsf{u}=[\mathsf{u}_{1},\mathsf{u}_{2},\mathsf{u}_{3}]^{\top}$ represents the torque input, and with $J_{1}=200,J_{2}=200,J_{3}=100$ denote the principal moments of inertia. $\tau=0.01$ is the sampling time. We consider the state space of the system to be ${\mathbf{X}}=[-0.25,0.25]\times[-0.25,0.25]\times[-0.25,0.25]$. Moreover, we consider the external input set to be bounded within ${\mathbf{W}}=[-1,1]$. Note that the internal input is of 3 dimensions while the external input is of a single dimension, i.e., ${\mathbf{U}}\in{\mathbb{R}}^{3}$, but $\mathbf{W}$$\in{\mathbb{R}}$. Also, we consider the model to be unknown. However, the Lipschitz constants $\mathcal{L}_{x}=1,\mathcal{L}_{u}=0.01$ are known for the system.
We fix the training hyper-parameters as $\epsilon=0.0125,\mathcal{L}_{L}=1,\mathcal{L}_{C}=40,{k=[0.00001,0.5,0.001,0.01],\Gamma=[2,2,2,2]}$. So, the Lipschitz constant according to Theorem 3.2 is $3.651$. We fix the structure of $V_{\theta,b}$ as $l_{v}=1,h_{v}^{1}=60$ and $g_{\bar{\theta},\bar{b}}$ as $l_{c}=1,h_{c}^{1}=40$. The activation function for both $\delta$-ISS-CLF and the controller is considered to be ‘ReLU’. The training algorithm converges to obtain $\delta$-ISS-CLF $V_{\theta,b}$ along with $\eta=-0.0460$. Hence, $\eta+\mathcal{L}\epsilon=-0.0460+3.651\times 0.0125=-0.00036$, therefore, using Theorem 3.5, we can guarantee that the obtained $\delta$-ISS-CLF $V_{\theta,b}$ is valid and that the closed-loop system is assured to be incrementally input-to-state stable under the influence of the controller $g_{\bar{\theta},\bar{b}}$.
The successful runs of the algorithm including the dataset generation time have an average convergence time of 5.5 hours. Note that the computational complexity grows rapidly with respect to system dimension as proposed in [32].
@@ 4.5 Discussion
The controller synthesis technique proposed in this work for achieving incremental input-to-state stability in closed-loop systems offers several advantages over existing methods. As the current technique relies on synthesizing the controller using data of the system, hence designing the controller does not require the system to be either fully actuated or affine structure as done in previous conventional techniques [18, 19]. This neural network based controller synthesis is able to make underactuated systems incrementally stable without the knowledge of dynamics. For instance, our neural network-based controller successfully ensures incremental stability in the one-link manipulator and jet-engine model without requiring detailed dynamic equations [42]. The approach is also reliable for the case of nonaffine systems as well, as seen in the scalar example.
Also, from the final case study, although the system has three independent control inputs, the neural controller can generate them using the system’s state and a single external input $w\in{\mathbf{W}}\subset{\mathbb{R}}$, so the closed-loop system becomes $\delta$-ISS with respect to a scalar exogenous input, which significantly simplifies control implementation. Now, as seen from [4], the computation complexity for the abstraction of an incrementally stable system lies in the dimension of the input to the system. As we have made the system incrementally stable with one input only, the computational complexity of the abstraction will reduce drastically once the controller is trained. This is helpful in finding the scalable symbolic abstraction of the system and, thus, will enable the efficient design of another formally verified controller to perform other specifications for the closed-loop system.
@@ 5 Conclusion and Future Work
In this paper, we presented a new training approach for unknown discrete-time systems that simultaneously synthesizes a controller and a verifiably correct incremental input-to-state stable control Lyapunov function, both parameterized as neural networks. The notion of incremental input-to-state state control Lyapunov function was introduced for the very first time. We highlighted the crucial challenges that are faced while solving an ROP to ensure the existence of $\delta$-ISS-CLF and thereby mapped the problem into SOP. Consequently, we presented a validity condition, by virtue of which solving SOP ensures satisfying the corresponding ROP, necessitating the existence of an appropriate $\delta$-ISS-CLF. We illustrated various case studies in support of our proposed claims. A promising future direction is to extend the framework to construct $\delta$-ISS-CLF for continuous-time systems. Moreover, we plan to extend these results to guarantee incremental ISS for different systems in stochastic settings where stochasticity might occur due to measurement noise and large-scale interconnected systems.
Appendix A: Proof of Theorem 2.7
The proof follows a similar approach of [12, Theorem 8]. Let the closed-loop system admit a $\delta$-ISS-CLF satisfying the conditions of Definition 2.6. The lower bound (i) and inequality (ii) imply: $\forall x,\hat{x}\in{\mathbf{X}},w,\hat{w}\in{\mathbf{W}}$:
[…]
@@
We denote $id$ as an identity function, i.e. $id(s):=s$ for all $s\in{\mathbb{R}}_{0}^{+}$. Assume $\chi$ be any class $\mathcal{K}_{\infty}$ such that $id-\chi\in\mathcal{K}_{\infty}$. Without loss of generality, we assume $id-\alpha\in\mathcal{K}$ [43, Lemma B.1].
Let $^{\mathsf{w}},\hat{{\mathsf{w}}}$ denote two input sequences such that ${\mathsf{w}},\hat{{\mathsf{w}}}:{\mathbb{N}}_{0}\rightarrow{\mathbf{W}}$. Now, define the set :
[…]
@@
Proof:
Let $(x,\hat{x})\in\mathbb{S}$. Also, ${\mathsf{w}}(0)=w,\hat{{\mathsf{w}}}(0)=\hat{w}\in{\mathbf{W}}$ be the first elements of $^{\mathsf{w}},\hat{{\mathsf{w}}}$ respectively. Note that the set $\mathbf{X}$} is compact and hence under the controller $g$, it is control forward invariant. Therefore, the terms $f(x,g(x,w))$ and $f(\hat{x},g(\hat{x},\hat{w}))\in{\mathbf{X}}$.
[…]
Since $(id-\chi)\in\mathcal{K}_{\infty}$, hence $(id-\chi)\circ\chi^{-1}\circ\sigma(\lVert{\mathsf{w}}-\hat{{\mathsf{w}}}\rVert)\leq 0$. So, essentially,
[…]
which implies that $(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))\in\mathbb{S}$. Hence, we can say that the set $\mathbb{S}$ is forward invariant. $\square$
@@
[…]
@@
[…]
$).\displaystyle V(f(x,g(x,w)),f(\hat{x},g(\hat{x},\hat{w})))-V(x,\hat{x})\leq-\alpha(V(x,\hat{x}))+\chi\circ\alpha(V(x,\hat{x}))\leq(id-\chi)\circ\alpha(V(x,\hat{x}$)).
Since, $id-\chi\in\mathcal{K}_{\infty}$, using the comparison lemma [44, Lemma 4.3], we infer there exists some $\tilde{\beta}\in\mathcal{KL}$ so that,
$\alpha_{1}(|\mathsf{x}(k,x,g(x,\mathsf{w}(k)))-\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k)))|)$ $\leq V(\mathsf{x}(k,x,g(x,\mathsf{w}(k))),\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k))))$
@@
$|\mathsf{x}(k,x,g(x,\mathsf{w}(k)))-\mathsf{x}(k,\hat{x},g(\hat{x},\mathsf{\hat{w}}(k)))|\leq\alpha_{1}^{-1}\circ\tilde{\beta}(\alpha_{2}(|x-\hat{x}|),k)=:\bar{\beta}((|x-\hat{x}|),k),$ (22)
@@
Now, combining (21) and (22), we obtain
[…]
@@
$\text{(c)}V_{\theta,b}(f(x,g_{\bar{\theta},\bar{b}}(x,w)),f(\hat{x},g_{\bar{\theta},\bar{b}}(\hat{x},\hat{w})))-V_{\theta,b}(x,\hat{x})+k_{3}|x-\hat{x}|^{\gamma_{3}}-k_{w}|w-\hat{w}|^{\gamma_{w}}$
@@
$\quad{+}V_{\theta,b}(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})),f(x_{r},g_{\bar{\theta},\bar{b}}(x_{r},w_{r})))+V_{\theta,b}(x_{q},x_{r})-V_{\theta,b}(x_{q},x_{r})+k_{3}|x-\hat{x}|^{\gamma_{3}}$
@@
$\leq\eta_{S}^{*}+\big(\sqrt{2}\mathcal{L}_{L}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+1)+2(\mathcal{L}_{3}+\mathcal{L}_{w})\big)\varepsilon\leq\eta_{S}^{*}+\mathcal{L}\varepsilon\leq 0.$
$\text{(d)}h(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-h(x)$
$=h(f(x,g_{\bar{\theta},\bar{b}}(x,w)))-h(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))+h(f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))-h(x)+h(x_{q})-h(x_{q})$
$\leq\mathcal{L}_{h}(f(x,g_{\bar{\theta},\bar{b}}(x,w))-f(x_{q},g_{\bar{\theta},\bar{b}}(x_{q},w_{q})))+\eta_{S}^{*}+\mathcal{L}_{h}\varepsilon$
$\leq\mathcal{L}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C})\varepsilon+\eta_{S}^{*}+\mathcal{L}_{h}\varepsilon\leq\eta_{S}^{*}+\mathcal{L}_{h}(\mathcal{L}_{x}+\sqrt{2}\mathcal{L}_{u}\mathcal{L}_{C}+1)\varepsilon\leq\eta_{S}^{*}+\mathcal{L}\varepsilon\leq 0.$
Therefore, if the condition (9) is satisfied, the neural $\delta$-ISS-CLF will ensure the system to be $\delta$-ISS under the controller.